OpenStudy (anonymous):

Please help me with the quadratic equation problem! A manufacturing crew needs to assemble 1000 boxes per day, divided equally among the workers. One day three workers call in sick, and the remaining members each need to assemble 75 more boxes than usual. How many workers are on the manufacturing crew?

OpenStudy (anonymous):

N= number workers b = boxes per worker N b = 1000 and (N-3)(b+75) = 1000 so Nb +75 N -3b - 225 = 1000 1000 + 75N - 3b = 1225 75N-3b = 225 multiply by N 75 N^2 -3Nb -225N =0 75 N^2 -225 N -3000=0 N = [225 +- sqrt(225^2 + (4)(75)(3000)]/150 N = [225 +- 975]/150 = 8 ignore negative result Nb = 1000 b = 1000/8 = 125 (8-3)(125+75) = 5(200) =1000 checks